Write an expression for semi-empirical formula for binding energy of nuclei given
by Weizsacker and discuss its each term. Using this formula, calculate the binding
energy per nucleon for
235 U.
1
Expert's answer
2020-04-09T09:15:53-0400

The semi-empirical formula for binding energy of nuclei given by Weizsacker is written below:


"B_E=a_vA-a_sA^{2/3}-a_c\frac{Z(Z-1)}{A^{1/3}}-a_a\frac{(N-Z)^2}{A}+a_p\frac{1}{A^{1/2}},"

where

"a_v=15.8\text{ MeV}\\
a_s=18.3\text{ MeV}\\
a_c=0.71\text{ MeV}\\
a_a=23.2\text{ MeV}\\
a_p=\pm11.2\text{ MeV}\\"

are best-fit parameters.,

"A" - nucleon number, "N" - neutron number, "Z -" proton number.

The first term with subscript v is volume term.

The second term with subscript s is surface term.

The third term with subscript c is coulomb term.

The fourth term with subscript a is asymmetry term.

The fifth term with subscript p is pairing term - defines the effect of spin-coupling. It's positive with Z, N, A even, zero with A odd, negative with Z, N odd and A even.

Therefore, for 235U this will give


"B_E=1796.1\text{ MeV}."